OpenStudy (rebeccapink6398):

Kyle paid $96 in interest to borrow $7200 for 4 months. What annual interest rate did Kyle pay? A. 1.33% B. 4% C. 5.33% D. 16%

OpenStudy (anonymous):

A

OpenStudy (anonymous):

As it is not specified in question if its being calculated using simple interest or compound interest I used compound interestone Using formula for compound interest : A=P(1+r/100)^n where A is amount,P is principal,r is rate and n is time So Interest charged,I =A-P or, I=P{[(1+(r/100))^n]-1} or, 96 =7200{[1+(r/4*100)^4]-1} or,96/7200=[1+(r/400)]^4]-1 or,7296/7200=[1+(r/400)]^4 or,\[\sqrt[4]{1.01333}\]=1+(r/400) or,1.0032=1+(r/400) or,r=1.327%~ 1.33% Option A If I would have used simple interest formula, my answer will be 1.2%